# VScan: Rethinking Visual Token Reduction for Efficient Large Vision-Language Models

### v2

@@ Abstract
Recent Large Vision-Language Models (LVLMs) have advanced multi-modal understanding by incorporating finer-grained visual perception and encoding. However, such methods incur significant computational costs due to longer visual token sequences, posing challenges for real-time deployment. To mitigate this, prior studies have explored pruning unimportant visual tokens either at the output layer of the visual encoder or at the early layers of the language model. In this work, we revisit these design choices and reassess their effectiveness through comprehensive empirical studies of how visual tokens are processed throughout the visual encoding and language decoding stages. Guided by these insights, we propose VScan, a two-stage visual token reduction framework that addresses token redundancy by: (1) integrating complementary global and local scans with token merging during visual encoding, and (2) introducing pruning at intermediate layers of the language model. Extensive experimental results across four LVLMs validate the effectiveness of VScan in accelerating inference and demonstrate its superior performance over current state-of-the-arts on sixteen benchmarks. Notably, when applied to LLaVA-NeXT-7B, VScan achieves a 2.91$\times$ speedup in prefilling and a 10$\times$ reduction in FLOPs, while retaining 95.4% of the original performance. Code is available at https://github.com/Tencent/SelfEvolvingAgent/tree/main/VScan.
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have emerged as a transformative advancement in multi-modal learning, achieving remarkable proficiency across a broad range of vision-language tasks [41; 34; 33; 54]. Recent advances in LVLMs [40; 32; 28; 6; 37; 46; 45] further enhance their capacity to process high-resolution images and multi-image/video inputs, enabling fine-grained perception in tasks such as video question answering [19; 59; 70], multi-image understanding [21; 29], and referential grounding [30; 47]. However, processing such rich visual inputs necessitates a substantial increase in the number of visual tokens, which often far exceeds the number of text tokens [37; 32]. For instance, LLaVA-NeXT [40] encodes up to 2,880 visual tokens for high-resolution images, while Qwen-2.5-VL [6] can process up to 16,384 tokens for multi-image or video inputs—orders of magnitude higher than typical text-only sequences. This leads to significantly longer input sequences and, due to the quadratic complexity of self-attention [56], incurs substantial computational and memory overhead, thereby limiting the scalability and real-time deployment of LVLMs in practical applications [12; 62].
Figure 1: Comparison of our VScan with existing text-agnostic approaches (e.g., VisionZip [62]) and text-aware approaches (e.g., FastV [12]). VScan is a two-stage training-free visual token reduction framework that can be seamlessly applied to various LVLM architectures, delivering significant acceleration in inference with minimal performance loss.
Recognizing that not all visual tokens contribute meaningfully to the final LVLM response, recent works [12; 60; 68] have proposed visual token reduction techniques aimed at improving computational efficiency by pruning visually redundant or textually irrelevant tokens. These methods generally fall into two categories: (1) Text-agnostic pruning approaches [68; 62; 57; 58] (Figure 1(a)), which prune visually redundant tokens based on their significance and uniqueness during the visual encoding stage, typically leveraging self-attention or [CLS] attention from the output layer of the visual encoder [36]; and (2) Text-aware pruning approaches [69; 60; 63] (Figure 1(b)), which selectively remove tokens with low relevance to the text query during the early layers of language decoding stage to preserve task-specific information while reducing computation. While these approaches have shown promising results, their performance is often constrained by their single-stage design and the lack of a systematic understanding of how visual tokens are processed and utilized throughout the entire LVLM pipeline.
@@
Building on these insights, we introduce VScan, a two-stage training-free visual token reduction framework that enhances the efficiency of LVLMs by progressively pruning uninformative tokens during both visual encoding and language decoding stages, as shown in Figure 1(c). In the visual encoding stage, VScan employs a complementary global-local scan strategy to retain semantically important and spatially diverse tokens, followed by token merging to preserve comprehensive visual information. In the LLM decoding stage, VScan introduces middle layer pruning to further eliminate visual tokens with low relevance to the text query, while maintaining essential cross-modal interactions to minimize disruption to final task performance. Notably, VScan can be seamlessly integrated into diverse open-sourced LVLM architectures and is fully compatible with FlashAttention [18; 17].
We comprehensively evaluate the effectiveness of VScan on LLaVA-1.5 [39], LLaVA-NeXT [40], Qwen-2.5-VL [6], and Video-LLaVA [37] across sixteen image and video understanding benchmarks. Extensive experimental results demonstrate VScan’s generalizable effectiveness across diverse LVLM architectures and LLM scales, highlighting its advantageous performance-efficiency trade-off. Specifically, VScan achieves a 1.77×1.$7$ speedup on LLaVA-1.5-7B and a 2.91×2.91$\times$2.91 × speedup on LLaVA-NeXT-7B during prefilling, while retaining 96.7% and 95.4% of the original performance, respectively.
@@ 2 Related Work
Efficient Large Vision-Language Models. Building on powerful auto-regressive LLMs [1; 55; 15], recent LVLMs typically adopt an encoder-projector-decoder architecture, where visual inputs are encoded into tokens and jointly processed with language sequences [41; 37; 5; 13; 54; 11]. However, as image resolution increases or the input scales to multi-image/video, the number of visual tokens grows proportionally, leading to a quadratic increase in computation cost and runtime due to the self-attention mechanism [56; 8; 14; 65], which limits the scalability of LVLMs in real-world applications [7; 12; 48; 31; 67; 66]. To mitigate this issue, several LVLMs introduced specialized modules to enhance efficiency—such as the Q-Former in InstructBLIP [16] and the perceiver resampler [26] in OpenFlamingo [2]—that distill dense visual inputs into a compact set of features before LLM decoding. Orthogonal to these architectural strategies, FlashAttention [18; 17] has emerged as a widely adopted, hardware-aware optimization that accelerates attention computation by minimizing redundant memory access, offering substantial speedups without compromising performance.
Vision Token Reduction in LVLMs. Another line of work aims to improve model efficiency on the sequence dimension—pioneering works such as ToMe [7] and FastV [12] have explored strategies like visual token merging and text-guided pruning to improve the efficiency of LVLMs. Building on these advances, subsequent approaches can be broadly divided into two main categories: (1) Text-agnostic pruning approaches [51; 4; 68; 62; 57; 58], which identify and remove redundant or uninformative visual tokens during the visual encoding stage. For instance, VisionZip [62] selects dominant tokens based on [CLS] attention scores, while FOLDER [57] introduces token merging with reduction overflow in the final blocks of the visual encoder. (2) Text-aware pruning approaches [69; 60; 42; 63; 53], which aim to remove visual tokens that are irrelevant to the text query during the LLM decoding stage. For instance, SparseVLM [69] proposes an iterative sparsification strategy that selects visual-relevant text tokens to rate the significance of vision tokens, and PyramidDrop [60] performs progressive pruning at multiple decoding layers to balance efficiency and context preservation. In this work, we present a comprehensive analysis of how LVLMs process visual tokens during both the visual encoding and language decoding stages, and propose a corresponding two-stage approach, VScan, to effectively improve the inference efficiency of LVLMs while maintaining robust performance.
@@ 3 Empirical Analysis
Rethinking Visual Redundancy Reduction. To address visual redundancy in token representations, recent studies [68; 62] have proposed text-agnostic approaches that retain visual tokens with high [CLS] attention at the output layer of the ViT-based visual encoder. However, this raises a critical question: Is relying solely on output [CLS] attention truly sufficient to capture all task-relevant visual information? Upon closer examination, we identify a clear yet often overlooked limitation of these approaches: they tend to favor tokens corresponding to visually salient objects, while aggressively discarding background visual details that may carry essential semantic information. As illustrated in Figure 2 (Left), output [CLS] attention is incorrectly directed to the wall and person, ignoring the actual targets—the pan and leather bag—leading to incorrect model responses.
@@
Figure 2: Empirical study on visual redundancy reduction. (Left) We illustrate two failure cases where relying solely on the output [CLS] attention leads to incorrect predictions. For comparison, we include reference token selections from CLIP-ViT-L-336px [50], following Gandelsman et al. [22], which highlight regions of interest relevant to the text query. (Right) We visualize the [CLS] attention maps and self-attention maps of representative tokens (e.g., #536: ground, #234: person) across different visual encoding layers, illustrating how attention patterns evolve from localized focus in shallow layers to broader global context in deeper layers.
Rethinking Textual Irrelevance Reduction. While prior studies [12; 69; 38] have proposed effective text-aware approaches for pruning visual tokens at early layers during LLM decoding, a critical question remains: Are early layers the optimal stage for pruning visual tokens to minimize their impact on the model’s final response? To investigate this, we conduct three empirical studies on POPE [35] and GQA [25], analyzing how the model’s knowledge and predictions evolve during the decoding process:
Figure 3: (Left) Study 1: Distribution of retained tokens at a 50% reduction rate in layers 2, 8, and 16 of LLaVA-1.5-7B on POPE [35]; (Right) Study 2: Sum of visual attention across different attention heads and LLM layers using LLaVA-1.5-7B and Qwen-2.5-VL-7B on POPE [35].
  • Study 1: How does position bias in token selection evolve across LLM layers? Specifically, we visualize the distribution of the retained tokens selected by the attention score of the last instruction token [12] across LLM layers using LLaVA-1.5-7B. As shown in Figure 3 (Left), early layers (e.g., layers 2 and 8) tend to select tokens at the bottom of the image, reflecting an inherent LLM position bias, as the last instruction token primarily attends to nearby tokens and focuses on local context [56; 49], and flattened visual tokens from the bottom of the image are positioned closest to the instruction tokens in the sequence. As the LLM layers deepen, this undesirable position bias diminishes and the focus shifts toward the center of the image, which is more intuitive since the center of the image typically carries the most informative and task-relevant features [3; 9].
@@
  • Study 3: At which LLM layer do next-token predictions begin to converge? In Figure 4, we provide an interpretation of the hidden states across different LLM layers in LLaVA-1.5-7B. Specifically, we feed the hidden states from each LLM decoding layer into the final linear projection layer to obtain vocabulary logits and intermediate next-token predictions. We observe that in more challenging open-ended tasks like GQA, the next-token predictions stabilize around LLM layer 20, whereas in simpler yes/no tasks such as POPE, the predictions converge earlier, around LLM layer 16.
@@
Figure 4: Study 3: Visualization of next-token predictions derived from the output hidden states of each LLM layer using LLaVA-1.5-7B. Darker colors indicate higher prediction confidence.
@@ 4.1 Reducing Visual Redundancy via Complementary Global and Local Scans
Global Scan. Given that the final layers of visual encoders capture global information, we follow recent works [68; 62] to select global tokens that receive the most attention from the [CLS] token $x_{[\mathtt{CLS}]}$ in the output layer (e.g., the penultimate layer in LLaVA-1.5 [39]). Specifically, the [CLS] attention computation for each attention head can be represented by
@@
Here, $\tau$ is a soft threshold based on a top percentile of attention scores, set to retain a target number of tokens. Note that for LVLMs without a [CLS] token (e.g., Qwen-2.5-VL [6]), we can similarly select the tokens using self-attention, i.e., the average attention each visual token receives from others.
@@
Token Merging. To alleviate information loss, we introduce a similarity-based token merging strategy that merges unselected visual tokens with their most similar selected counterparts. Specifically, for each unselected token $x_{V}^{u}$, we identify its most similar selected token $x_{V}^{s}\in\mathbf{x}_{V}^{\mathtt{selected}}$ based on the highest cosine similarity. Once all unselected tokens are assigned to their closest selected tokens, we apply average merging [7] within each group to obtain the final merged representation $\mathbf{x}_{V}^{\mathtt{merged}}$.
@@ 4.2 Reducing Textual Irrelevance via Middle Layer Pruning
We similarly average the attention scores across different attention heads and select $R_{2}$textually relevant tokens with the highest average text attention. This allows us to retain a set of visual tokens that are both visually significant and textually relevant, contributing the most to an accurate response.
@@
Empirical Validation. We conduct a comparative analysis on the GQA benchmark using LLaVA-1.5-7B [39] to examine the effect of pruning tokens at different LLM layers, while keeping the average reduction rate consistent across settings. To better highlight the impact of pruning depth, we first apply a global scan to reduce the visual tokens to 288/144 (i.e., $R_{1}$ = 50%/25%), and then perform pruning at various LLM layers to reach an average retention rate of 75% during the LLM decoding stage. As shown in Table 1, pruning at middle LLM layers (e.g., layers 16 or 20) yields the best performance, whereas pruning at earlier layers (e.g., layer 2) leads to up to a 1.9% drop in accuracy. These results align with our empirical insights in Section 3 and validate the effectiveness of pruning at middle layers to remove textual irrelevance in LVLMs.
Remarks on FlashAttention. Our proposed VScan is compatible with FlashAttention [17; 18], as we recompute the attention scores for the last instruction token using vanilla attention calculation [56] outside the standard LLM layers. A detailed efficiency analysis can be found in Section 5.4.
Table 2: Performance comparisons on LLaVA-1.5-7B [39] across 9 image understanding benchmarks. The best results in each setting are bolded, and the second-best are underlined.
[…]
Upper Bound, 576 Tokens (100%), 3.817 TFLOPs
LLaVA-1.5-7B [39] CVPR’24 61.9 64.7 58.1 1862 85.9 69.5 78.5 58.2 50.0 100.0%
Retain 192 Tokens in Average ($\downarrow$ 66.7%), $\sim$1.253 TFLOPs
ToMe [7] ICLR’23 54.3 60.5 - 1563 72.4 65.2 68.0 52.1 - 88.5%
FastV [12] ECCV’24 52.7 61.2 57.0 1612 64.8 67.3 67.1 52.5 50.8 90.4%
SparseVLM [69] ICML’25 57.6 62.5 53.7 1721 83.6 69.1 75.6 56.1 50.5 96.1%
PyramidDrop [60] CVPR’25 57.3 63.3 56.8 1797 82.3 69.0 75.1 56.5 51.1 97.2%
VisionZip [62] CVPR’25 59.3 63.0 - 1783 85.3 68.9 77.4 57.3 - 97.8%
VScan (Ours) - 60.6 63.9 57.4 1806 86.2 68.6 77.8 57.7 50.4 99.0%
Retain 128 Tokens in Average ($\downarrow$ 77.8%), $\sim$0.833 TFLOPs
ToMe [7] ICLR’23 52.4 53.3 - 1343 62.8 59.6 63.0 49.1 - 80.4%
FastV [12] ECCV’24 49.6 56.1 56.4 1490 59.6 60.2 61.8 50.6 51.3 85.4%
SparseVLM [69] ICML’25 56.0 60.0 51.1 1696 80.5 67.1 73.8 54.9 51.4 93.7%
PyramidDrop [60] CVPR’25 57.1 61.6 56.6 1761 82.3 68.4 72.9 56.6 51.0 96.2%
VisionZip [62] CVPR’25 57.6 62.0 - 1763 83.2 68.9 75.6 56.8 - 96.2%
VScan (Ours) - 59.8 63.0 58.0 1792 86.1 68.9 77.1 57.3 51.7 98.8%
Retain 64 Tokens in Average ($\downarrow$ 88.9%), $\sim$0.415 TFLOPs
ToMe [7] ICLR’23 48.6 43.7 - 1138 52.5 50.0 57.1 45.3 - 70.1%
FastV [12] ECCV’24 46.1 48.0 52.7 1256 48.0 51.1 55.0 47.8 50.8 76.7%
SparseVLM [69] ICML’25 52.7 56.2 46.1 1505 75.1 62.2 68.2 51.8 50.1 87.2%
PyramidDrop [60] CVPR’25 47.5 58.8 50.5 1561 55.9 69.2 69.2 50.6 50.7 86.6%
VisionZip [62] CVPR’25 55.1 60.1 - 1690 77.0 69.0 72.4 55.5 - 92.7%
VScan (Ours) - 58.3 62.1 55.7 1698 85.0 69.1 75.4 55.6 51.8 96.7%

### v3

@@ Abstract
[…]
@@ 1 Introduction
Large Vision-Language Models (LVLMs) have emerged as a transformative advancement in multi-modal learning, achieving remarkable proficiency across a broad range of vision-language tasks Liu et al. [2023]; Li et al. [2024; 2023a]; Team et al. [2024]. Recent advances in LVLMs Liu et al. [2024b]; Li et al. [2025]; Jia et al. [2024]; Bai et al. [2025]; Lin et al. [2024]; Maaz et al. [2024]; Ma et al. [2024] further enhance their capacity to process high-resolution images and multi-image/video inputs, enabling fine-grained perception in tasks such as video question answering Fang et al. [2024]; Wu et al. [2024]; Song et al. [2025], multi-image understanding Fu et al. [2024]; Jiang et al. [2024], and referential grounding Kazemzadeh et al. [2014]; Mao et al. [2016]. However, processing such rich visual inputs necessitates a substantial increase in the number of visual tokens, which often far exceeds the number of text tokens Lin et al. [2024]; Li et al. [2025]. For instance, LLaVA-NeXT Liu et al. [2024b] encodes up to 2,880 visual tokens for high-resolution images, while Qwen-2.5-VL Bai et al. [2025] can process up to 16,384 tokens for multi-image or video inputs—orders of magnitude higher than typical text-only sequences. This leads to significantly longer input sequences and, due to the quadratic complexity of self-attention Vaswani et al. [2017], incurs substantial computational and memory overhead, thereby limiting the scalability and real-time deployment of LVLMs in practical applications Chen et al. [2024a]; Yang et al. [2025].
Figure 1: Comparison of our VScan with representative text-agnostic approaches (e.g., VisionZip Yang et al. [2025]) and text-aware approaches (e.g., FastV Chen et al. [2024a]). In this work, we introduce VScan, a two-stage, training-free visual token reduction framework that can be seamlessly applied to various open-sourced LVLM architectures, delivering significant acceleration in inference with minimal performance loss.
Recognizing that not all visual tokens contribute meaningfully to the final LVLM response, recent works Chen et al. [2024a]; Xing et al. [2025]; Zhang et al. [2024] have proposed visual token reduction techniques aimed at improving computational efficiency by pruning visually redundant or textually irrelevant tokens. These methods generally fall into two categories: (1) Text-agnostic pruning approaches Yang et al. [2025]; Wang et al. [2025]; Wen et al. [2025] (Figure 1(a)), which prune visually redundant tokens based on their significance and uniqueness during the visual encoding stage, typically leveraging self-attention or [CLS] attention from the output layer of the visual encoder; and (2) Text-aware pruning approaches Zhang et al. [2025c]; Xing et al. [2025]; Ye et al. [2025] (Figure 1(b)), which selectively remove tokens with low relevance to the text query during the early layers of language decoding stage to preserve task-specific information while reducing computation. While these approaches have shown promising results, their performance is often constrained by their single-stage design and the lack of a systematic understanding of how visual tokens are processed and utilized throughout the entire LVLM pipeline.
@@
Building on these insights, we introduce VScan, a two-stage, training-free visual token reduction framework that enhances the efficiency of LVLMs by progressively pruning uninformative tokens during both visual encoding and language decoding stages, as shown in Figure 1(c). In the visual encoding stage, VScan employs a complementary global-local scan strategy to retain semantically important and spatially diverse tokens, followed by token merging to preserve comprehensive visual information. In the LLM decoding stage, VScan introduces middle layer pruning to further eliminate visual tokens with low relevance to the text query, while maintaining essential cross-modal interactions to minimize disruption to final task performance. Notably, VScan can be seamlessly integrated into diverse open-sourced LVLM architectures and is fully compatible with FlashAttention Dao et al. [2022]; Dao [2024], making it both practical and broadly applicable to real-world applications.
We comprehensively evaluate the effectiveness of VScan on LLaVA-1.5 Liu et al. [2024a], LLaVA-NeXT Liu et al. [2024b], Qwen-2.5-VL Bai et al. [2025], and Video-LLaVA Lin et al. [2024] across sixteen image and video understanding benchmarks. Extensive experimental results demonstrate VScan’s generalizable effectiveness across diverse LVLM architectures and LLM scales, highlighting its advantageous performance-efficiency trade-off. Specifically, VScan achieves a $1.77\times$ speedup on LLaVA-1.5-7B and a $2.91\times$ speedup on LLaVA-NeXT-7B during prefilling, while retaining 96.7% and 95.4% of the original performance, respectively.
@@ 2 Related Work
Efficient Large Vision-Language Models. Building on powerful auto-regressive LLMs Touvron et al. [2023]; Chung et al. [2024], recent LVLMs typically adopt an encoder-projector-decoder architecture, where visual inputs are encoded into tokens and jointly processed with language sequences Liu et al. [2023]; Lin et al. [2024]; Bai et al. [2023]; Chen et al. [2024b]; Team et al. [2024]; Chen et al. [2023]. However, as image resolution increases or the input scales to multi-image/video, the number of visual tokens grows proportionally, leading to a quadratic increase in computation cost and runtime due to the self-attention mechanism Vaswani et al. [2017]; Brauwers and Frasincar [2021]; Child et al. [2019]; Zaheer et al. [2020], which limits the scalability of LVLMs in real-world applications Bolya et al. [2023]; Chen et al. [2024a]; Mehta and Rastegari [2022]; Kondratyuk et al. [2021]; Zhang et al. [2025b; a]; Chai et al. [2025]. To mitigate this issue, several LVLMs introduced specialized modules to enhance efficiency—such as the Q-Former in InstructBLIP Dai et al. [2023] and the perceiver resampler Jaegle et al. [2021] in OpenFlamingo Alayrac et al. [2022]—that distill dense visual inputs into a compact set of features before LLM decoding. Orthogonal to these architectural strategies, FlashAttention Dao et al. [2022]; Dao [2024] has emerged as a widely adopted, hardware-aware optimization that accelerates attention computation by minimizing redundant memory access, offering substantial speedups without compromising performance.
Vision Token Reduction in LVLMs. Another line of work aims to improve model efficiency on the sequence dimension—pioneering works such as ToMe Bolya et al. [2023] and FastV Chen et al. [2024a] have explored strategies like visual token merging and text-guided pruning to improve the efficiency of LVLMs. Building on these advances, subsequent approaches can be broadly divided into two main categories: (1) Text-agnostic pruning approaches Shang et al. [2024]; Arif et al. [2025]; Zhang et al. [2024]; Yang et al. [2025]; Wang et al. [2025]; Wen et al. [2025], which identify and remove redundant or uninformative visual tokens during the visual encoding stage. For instance, VisionZip Yang et al. [2025] selects dominant tokens based on [CLS] attention scores, while FOLDER Wang et al. [2025] introduces token merging with reduction overflow in the final blocks of the visual encoder. (2) Text-aware pruning approaches Zhang et al. [2025c]; Xing et al. [2025]; Liu et al. [2024c]; Ye et al. [2025]; Tan et al. [2025], which aim to remove visual tokens that are irrelevant to the text query during the LLM decoding stage. For instance, SparseVLM Zhang et al. [2025c] proposes an iterative sparsification strategy that selects visual-relevant text tokens to rate the significance of vision tokens, and PyramidDrop Xing et al. [2025] performs progressive pruning at multiple decoding layers to balance efficiency and context preservation. In this work, we present a comprehensive analysis of how LVLMs process visual tokens during both the visual encoding and language decoding stages, and propose a corresponding two-stage approach, VScan, to effectively improve the inference efficiency of LVLMs while maintaining robust performance.
@@ 3 Empirical Analysis
Rethinking Visual Redundancy Reduction. To address visual redundancy in token representations, recent studies Zhang et al. [2024]; Yang et al. [2025] have proposed text-agnostic approaches that retain visual tokens with high [CLS] attention at the output layer of the ViT-based visual encoder. While effective to some extent, this strategy raises an important question: Is relying solely on output [CLS] attention truly sufficient to capture all task-relevant visual information? Upon closer examination, we identify a clear yet often overlooked limitation of these approaches: they tend to favor tokens corresponding to visually salient objects, while aggressively discarding background visual details that may carry essential semantic information. As illustrated by the examples in Figure 2 (Left), output [CLS] attention is incorrectly directed to the wall and person, ignoring the actual targets—the pan and leather bag—leading to incorrect model responses.
@@
Figure 2: Empirical study on visual redundancy reduction. (Left) We illustrate two failure cases where relying solely on the output [CLS] attention leads to incorrect predictions. For comparison, we include reference token selections from CLIP-ViT-L-336px Radford et al. [2021], following Gandelsman et al. Gandelsman et al. [2024], which highlight regions of interest relevant to the text query. (Right) We visualize the [CLS] attention maps and self-attention maps of representative tokens (e.g., #536: ground, #234: person) across different encoding layers, illustrating how attention patterns evolve from localized focus in shallow layers to broader global context in deeper layers.
Rethinking Textual Irrelevance Reduction. While prior studies Chen et al. [2024a]; Zhang et al. [2025c]; Lin et al. [2025] have proposed effective text-aware approaches for pruning visual tokens at early layers during LLM decoding, a critical question remains: Are early layers the optimal stage for pruning visual tokens to minimize their impact on the model’s final response? To investigate this, we conduct three empirical studies on POPE Li et al. [2023b] and GQA Hudson and Manning [2019], analyzing how the model’s knowledge and predictions evolve during the decoding process:
Figure 3: (Left) Study 1: Distribution of retained tokens at a 50% reduction rate in layers 2, 8, and 16 of LLaVA-1.5-7B on POPE Li et al. [2023b]; (Right) Study 2: Sum of visual attention across different attention heads and LLM layers using LLaVA-1.5-7B and Qwen-2.5-VL-7B on POPE Li et al. [2023b].
Figure 4: Study 3: Visualization of next-token predictions derived from the output hidden states of each LLM layer using (a) LLaVA-1.5-7B; (b) LLaVA-NeXT-7B. Darker colors indicate higher prediction confidence.
  • Study 1: How does position bias in token selection evolve across LLM layers? Specifically, we visualize the distribution of the retained tokens selected by the attention score of the last instruction token Chen et al. [2024a] across LLM layers using LLaVA-1.5-7B. As shown in Figure 3 (Left), early layers (e.g., layers 2 and 8) tend to select tokens at the bottom of the image, reflecting an inherent LLM position bias, as the last instruction token primarily attends to nearby tokens and focuses on local context Vaswani et al. [2017]; Parmar et al. [2018], and flattened visual tokens from the bottom of the image are positioned closest to the instruction tokens in the sequence. As the LLM layers deepen, this undesirable position bias diminishes and the focus shifts toward the center of the image, which is more intuitive since the center of the image typically carries the most informative and task-relevant features Alexey et al. [2015]; Carion et al. [2020].
@@
  • Study 3: At which LLM layer do next-token predictions begin to converge? In Figure 4, we provide an interpretation of the hidden states across different LLM layers using LLaVA-1.5-7B and LLaVA-NeXT-7B. Specifically, we feed the hidden states from each LLM decoding layer into the final linear projection layer to obtain vocabulary logits and intermediate next-token predictions. We observe that in more challenging open-ended tasks like GQA, the next-token predictions stabilize around LLM layer 20, whereas in simpler yes/no tasks such as POPE, the predictions converge earlier, around LLM layer 16. Our findings indicate that early layers are still forming core cross-modal semantics, and pruning them risks disrupting essential grounding. In contrast, by the middle layers, next-token predictions have largely stabilized, meaning that these layers contribute diminishing semantic change. This directly motivates pruning in the middle-to-late layers rather than the early layers.
@@ 4.1 Reducing Visual Redundancy via Complementary Global and Local Scans
Global Scan. Given that the final layers of visual encoders capture global information, we follow recent works Zhang et al. [2024]; Yang et al. [2025] to select global tokens that receive the most attention from the [CLS] token $x_{[\mathtt{CLS}]}$ in the output layer (e.g., the penultimate layer in LLaVA-1.5 Liu et al. [2024a]). Specifically, the [CLS] attention computation for each attention head can be represented by
@@
Here, $\tau$ is a soft threshold based on a top percentile of attention scores, set to retain a target number of tokens. Note that for LVLMs without a [CLS] token (e.g., Qwen-2.5-VL Bai et al. [2025]), we can similarly select the tokens using self-attention, i.e., the average attention each visual token receives from others.
@@
Token Merging. To alleviate information loss, we introduce a similarity-based token merging strategy that merges unselected visual tokens with their most similar selected counterparts. Specifically, for each unselected token $x_{V}^{u}$, we identify its most similar selected token $x_{V}^{s}\in\mathbf{x}_{V}^{\mathtt{selected}}$ based on the highest cosine similarity. Once all unselected tokens are assigned to their closest selected tokens, we apply average merging Bolya et al. [2023] within each group to obtain the final merged representation $\mathbf{x}_{V}^{\mathtt{merged}}$. Specifically, for each selected token $\mathbf{x}_{V}^{s}$, we compute the average token representation by
$\mathbf{x}_{V}^{\mathtt{merged}}=\frac{1}{|\mathcal{U}^{s}|+1}\left(\sum_{u\in\mathcal{U}^{s}}x_{V}^{u}+\mathbf{x}_{V}^{s}\right),$ (4)
where $\mathcal{U}^{s}$ denotes the set of unselected tokens associated with the selected token $\mathbf{x}_{V}^{s}$, and $|\mathcal{U}^{s}|$ indicates the cardinality of this set.
@@ 4.2 Reducing Textual Irrelevance via Middle Layer Pruning
We similarly average the attention scores across different attention heads and select $R_{2}$% textually relevant tokens with the highest average text attention. This allows us to retain a set of visual tokens that are both visually significant and textually relevant, contributing the most to an accurate response.
@@
Empirical Validation. We conduct a comparative analysis on the GQA benchmark using LLaVA-1.5-7B Liu et al. [2024a] to examine the effect of pruning tokens at different LLM layers, while keeping the average reduction rate consistent across settings. To better highlight the impact of pruning depth, we first apply a global scan to reduce the visual tokens to 288/144 (i.e., $R_{1}$ = 50%/25%), and then perform pruning at various LLM layers to reach an average retention rate of 75% during the LLM decoding stage. As shown in Table 1, pruning at middle LLM layers (e.g., layers 16 or 20) yields the best performance, whereas pruning at earlier layers (e.g., layer 2) leads to up to a 1.9% drop in accuracy. These results align with our empirical insights in Section 3 and validate the effectiveness of pruning at middle layers to remove textual irrelevance in LVLMs.
Remarks on KV Cache and FlashAttention. Our proposed VScan is fully compatible with standard KV caching mechanisms, as pruning occurs before visual tokens are added to the cache. Consequently, the KV cache stores fewer entries while its structure and format remain unchanged. VScan is also compatible with FlashAttention Dao [2024]; Dao et al. [2022], as we recompute the attention scores for the last instruction token using vanilla attention calculation Vaswani et al. [2017] outside the standard LLM layers.